People Vw Updates Ready to Inst o restart to Chater C 0 Not Secure I ezto. m.tpx 0.6931122424484223,1548459013627 ncome of $6.900,000, total equity of 521,710,000, sales of $40,300,000, and Ames, Isc, has a current stock price of $45.00. For the past year, the oompany had net 5.4 milion shares of stock outstanding What are eamings per share (EPS)? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g, 32.16 Eamings per share $ 1.28 What is the price-eamings ratio? (De not round intermedjate calculations and round your answer to 2 decimal places,-g.32.16) 35 16 Smes What is the price-sales ratio? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g, 32.16) Price-sales What is the book value per share? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g.32.16 Book value per share What is the market to-book ratio? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g. 32.16. Market-to-book ratio Referances Book & Resources Worksheet Leaming Objective: 03-02 Compute and, more important interpret sene ㏄mmon . Please help, I don’t know how to do the last 3
Answer #1

Price to Sales Ratio = Price/( Sales/No of Shares ) = 45/(40,300,000/5,400,000) = 6.03

Book Value Per share = Total Equity/No of Shares = 21,710,000/5,400,000 = 4.02

Market To Book Ratio = Price Per Share/Book Value Per share = 45/4.02 =11.19

Please Discuss in case of Doubt

Best of Luck. God Bless
Please Rate Well